OpenStudy (anonymous):

Find a polynomial function P(x) of least possible degree, having real coefficients, with the zeros 2 and 3i. Let the leading coefficient be 1.

OpenStudy (deoxna):

Can the zeroes only be 2 and 3i, can there be more zeroes?

OpenStudy (anonymous):

The problem only listed 2 and 3i :/ Here is how I started the problem: 1(x-2)((x-(x-3i))((x-(x+3i))

OpenStudy (anonymous):

now just foil those.

OpenStudy (anonymous):

so 9x-18???

OpenStudy (anonymous):

i didnt do the math. but i dont see how you come out with a linear term with 3 x's in there. you should have a cubic function

OpenStudy (anonymous):

1(x-2)((x-(x-3i))((x-(x+3i)) -- In this expression you have subtracted x - (x - 3i) so the two x's cancel out. But it should be just x - 3i in place of x - (x - 3i) and the same for x - (x + 3i).

OpenStudy (anonymous):

yep, i didnt pay enough attention to catch that. good eye

OpenStudy (dumbcow):

here is simpler way to deal with imaginary zeros \[x = 3i\] \[x^2 = 9i^2\] \[x^2 = -9\] \[x^2 +9 = 0\] polynomial is: --> (x-2)(x^2 +9)

OpenStudy (anonymous):

thats interesting. it looks so simple. why isnt it taught that way? and what makes that legal to do out of curiosity?

OpenStudy (anonymous):

i guess. it just makes sense. because it is a zero of the polynomial.

OpenStudy (dumbcow):

right, it works because you are just working backwards start with solution then work back until right side is zero, then you have the factor x = 2 x-2 = 0

OpenStudy (anonymous):

i wish they taught it that way. makes doing the math a lot quicker

OpenStudy (dumbcow):

yep some teachers do

OpenStudy (anonymous):

x3−8x2+22x−20 is what I got for the polynomial :/ I broke it down to (x-2)(x-3-i)(x-3+i) and then multiplied.

OpenStudy (anonymous):

Multiply each term in the first polynomial by each term in the second polynomial

OpenStudy (dumbcow):

no try my method (x-2)(x^2 +9) = x^3 -2x^2 +9x -18

OpenStudy (anonymous):

okie. However I'm still a little unclear to where the x^2+9 came from...

OpenStudy (dumbcow):

can you not see my earlier post above?? i show steps

OpenStudy (anonymous):

so you put: x=3i because 3i is one of the zeros. and you squared it. X^2+9=3i Okay, I see how you got there. Thank you! i'm going to try your way on similar problems till I get it down :)

OpenStudy (dumbcow):

:) it works for when the zero is a sqrt too anytime the zero is an imaginary or sqrt they come in pairs so the factor is always a quadratic if have any other questions concerning this tag me

OpenStudy (anonymous):

awesome! made things much simpler